OpenStudy (anonymous):

Solve the following system of equations: x − 2y = 14 x + 3y = 9 (1, 12) (−1, −12) (12, −1) (12, 1)

OpenStudy (anonymous):

So you want to eliminate a component to make it easier to find the other one You know you can rearrange both of these equations for x x=14+2y x=9-3y So now you know 14+2y=9-3y put all your y's on one side 5y=-5 divide for y then substitute your 'y' value back in to your original equation to find 'x' hope this helps!

OpenStudy (texaschic101):

do you need help ?

OpenStudy (texaschic101):

sarahusher explained it very well

OpenStudy (texaschic101):

you could also multiply the 1st equation by -1. Then add the equations, ..you see how the x's cancel out, and you can solve for y. Once y is found, sub that back into either equation and solve for x.